OpenStudy (anonymous):

How many grams of N2 are in 5.6L of N2 at stp

OpenStudy (anonymous):

22.4 litres of a substance contain one mole of that substance. 1 mole=22.4 l Therefore... 5.6 l=(5.6/24) moles Now...we know that No of moles=given weight/weight of the substance 5.6/22.4=x/28. Where x is the grams of n2 present Solve for x